The rule of 70 states that an investment will double in 70/x years, where x is the annual rate of return rate.

Question 22 options:

a)

True

b)

False

John Stossel followed his "dart picked" portfolio for a year and found it beat 90% of the experts.

Question 23 options:

a)

True

b)

False

A mutual fund manager must demonstrate high ability to beat the market over a 1-year span.

Question 12 options:

a)

True

b)

False

The riskiest stocks are those that have a high positive correlation with the market as a whole.

Question 6 options:

a)

True

b)

False

In a market of 2,000 investors who each year flip a coin to predict market success or failure, how many investors will have been consistently right after 5 years? (Assume the coin tosses yield heads exactly 50% of the time.)

Question 25 options:

a)

31

b)

62

c)

500

d)

1,000
Answer #1

Answer 1: True

Answer 2: True

Answer 3 : False

Answer 4: True

Answer 5 a 31

First year 1000/2 = 500

II year 500/2 = 250

III year 250/2 = 125

IV year 125/2 = 62.5

v year 62.5 /2 - 31.25